OpenStudy (anonymous):

The switch for the turn table on 33rpm stereo system is turned off. The table makes 20 complete ratations before slowing to stop.Calculate it's angular acceleration.

OpenStudy (anonymous):

\[\alpha = {{\Delta \omega} \over {\Delta t}}\] 33 rpm = 33 rotation per minutes = 33 rotation / 60 seconds = x (you need to calculate this) Complete stop = 0 rad/s. So \[\Delta \omega\] = x - 0 rad/s = x \[\Delta t\] = 20/x seconds = t so \[\alpha\] = x/t (you need to compute this), it is in rad/s^2